With an initial velocity of 17 m/sec at a 70 degree angle, what is the maximum height of a canonball shot from a cannon?
Answer #1

Maximum height, h = u^2 (sin(theta))^2/2g

h = 17^2 x (sin70)^2/ 2 x 9.8 = 13.02 m

Comment in case any doubt please rate my answer ...